OpenStudy (anonymous):

Solve the following equations for all reals. sin^2 x - sin x = 0 sin 2x + cos x = 0

OpenStudy (anonymous):

For the first equation, factor out a sinx. sinx(sinx-1)=0 Set both factors equal to 0 So, find when sinx=0 and sinx=1 so x= 0, π, 2π, and π/2. Add a +2πn to each answer since there's an infinite number if it's asking for all reals. x= 0+2πn, π+2πn, π/2 + 2πn

OpenStudy (anonymous):

factor out a sin x huh? I dont know what you mean sinx * sinx -sinx =0?

OpenStudy (anonymous):

For the second equation, you have to know that sin2x = 2sinxcosx, so you have 2sinxcosx + cosx =0 Factor out a cosx. cosx(2sinx + 1) = 0 Set both factors equal to 0, so cosx=0 and sinx= -1/2 x= π/2 + 2πn, 3π/2 + 2πn x= 7π/6 + 2πn, 11π/6 + 2πn

OpenStudy (anonymous):

For the first equation, you have sin^2x, or sinx squared, - sinx. You need to factor out a sinx and you get sinx*(sinx-1)=0 (you can make sure the equations are equivalent by foiling. You also need to factor for the second equation.